# Enabling the Write-Back Page Cache with Strong Consistency in Distributed Userspace File Systems

### v2

@@ 6.4. Scalability (Q3)
Figure 8 illustrates DistFUSE’s scalability as the number of nodes increases from 2 to 16. The workload is 50%-50% random read-write with fio, and we set the contention level to 50%. Both DistFUSE and the baseline demonstrate near-linear scaling, with throughput increasing proportionally to the number of nodes, suggesting neither system suffers from significant coordination bottlenecks.
At smaller scales (2-4 nodes), DistFUSE demonstrates a significant performance advantage over the baseline, achieving approximately 18-21% higher throughput. This early advantage highlights that DistFUSE’s kernel-level lease management approach delivers benefits even in modest deployments common in departmental or small enterprise environments.
As the deployment scales to medium size (6-10 nodes), DistFUSE maintains a consistent performance lead of approximately 19-22% over the baseline. This steady advantage demonstrates that DistFUSE’s distributed coordination mechanisms scale efficiently without introducing significant overhead as the system grows. For instance, at 8 nodes, DistFUSE achieves 91 MB/s versus the baseline’s 76 MB/s, a performance difference that would be meaningful in production environments.
At larger scales (12-16 nodes), DistFUSE continues to outperform the baseline, though the relative advantage gradually diminishes from 14.8% at 12 nodes to 8.6% at 16 nodes. This trend suggests that while DistFUSE remains more efficient throughout the tested range, some aspects of its coordination mechanisms may benefit from further optimization for very large deployments.
Importantly, DistFUSE achieves this superior performance while maintaining strong consistency guarantees across all nodes and scales. The linear scaling pattern observed up to 16 nodes indicates that DistFUSE’s architecture successfully addresses the scalability requirements of modern cloud environments without compromising on consistency semantics.
These results validate DistFUSE’s design approach of offloading lease management to the kernel driver, demonstrating that it effectively balances performance, consistency, and scalability for distributed file system deployments of varying sizes.
@@ 7. Related Work
DistFUSE builds upon and extends prior work in distributed file systems, FUSE optimizations, and consistency mechanisms in distributed file systems. This section contextualizes our contributions within the broader research landscape.
@@ 7.1. FUSE-based Distributed File Systems
Unlike these systems, DistFUSE uniquely bridges the consistency-performance gap by enabling write-back caching with strong consistency through kernel-level lease management, avoiding the traditional tradeoff between performance and correctness.
@@ 7.2. FUSE Performance Optimizations
The Linux kernel has incorporated various FUSE optimizations, including the ability to splice read and write operations that move data between kernel buffers without copying to userspace(noa, 2010), significantly reducing memory copy overhead. Recent io_uring-based FUSE implementations(noa, 2024) further improve performance by avoiding core switching and using shared memory for data transfers. These approaches collectively address the fundamental bottlenecks in FUSE: frequent context switches, memory copies, and lock contention.
These approaches primarily optimize local FUSE performance, whereas DistFUSE addresses the fundamental distributed consistency challenge while leveraging these performance improvements. By offloading lease management to the kernel driver, DistFUSE reduces the performance impact of maintaining strong consistency in distributed environments.
7.3. Consistency Mechanism in Distributed File System
Distributed file systems employ various mechanisms to maintain consistency. Lease-based approaches, as pioneered by Gray and Cheriton(Gray and Cheriton, 1989), provide time-bound permissions that balance performance with consistency. GPFS (now Spectrum Scale)(Schmuck and Haskin, 2002) implements distributed locking for strong consistency but requires specialized hardware for optimal performance. Lustre(Schwan et al., 2003) employs a distributed lock manager to coordinate file access across nodes, though with different architectural tradeoffs than DistFUSE.
DistFUSE’s innovation lies in embedding lease management directly in the kernel’s FUSE driver while maintaining userspace flexibility. This approach differs fundamentally from previous systems that either implement consistency mechanisms entirely in userspace (sacrificing performance) or require extensive kernel modifications (sacrificing flexibility and isolation).
@@ 8. Conclusions
This paper presents DistFUSE, a FUSE-based distributed file system that reconciles write-back kernel page caching with strong consistency for cloud environments. DistFUSE eliminates the traditional trade-off between performance and correctness by offloading distributed lease management to the kernel driver, enabling coordinated access to cached data across nodes while retaining FUSE’s userspace flexibility. Our experiments demonstrate DistFUSE’s efficacy: it achieves up to 68.0% higher throughput and 40.4% lower latency than write-through baselines for write-heavy workloads and near-linear scalability to 16 nodes while maintaining consistency under contention. Application benchmarks show 36.2% throughput improvements and 30.0% latency reduction in contended Fileserver workloads, validating its practicality for real-world deployments. By embedding lease coordination in the kernel, DistFUSE demonstrates that distributed file systems can deliver both high performance and strong consistency, advancing the viability of FUSE-based solutions for cloud-scale workloads.

### v3

@@ 6.4. Scalability (Q3)
Figure 8 illustrates DFUSE’s scalability as the number of nodes (i.e., the count of client machines) increases from 2 to 16. The workload is 50%-50% random read-write with fio, and we set the contention level to 50%. Both DFUSE and the baseline demonstrate near-linear scaling, with throughput increasing proportionally to the number of nodes, suggesting neither system suffers from significant coordination bottlenecks.
At smaller scales (2-4 nodes), DFUSE demonstrates a significant performance advantage over the baseline, achieving approximately 18-21% higher throughput. This early advantage highlights that DFUSE’s kernel-level lease management approach delivers benefits even in modest deployments common in departmental or small enterprise environments.
As the deployment scales to medium size (6-10 nodes), DFUSE maintains a consistent performance lead of approximately 19-22% over the baseline. This steady advantage demonstrates that DFUSE’s distributed coordination mechanisms scale efficiently without introducing significant overhead as the system grows. For instance, at 8 nodes, DFUSE achieves 91 MB/s versus the baseline’s 76 MB/s, a performance difference that would be meaningful in production environments.
At larger scales (12-16 nodes), DFUSE continues to outperform the baseline, though the relative advantage gradually diminishes from 14.8% at 12 nodes to 8.6% at 16 nodes. This trend suggests that while DFUSE remains more efficient throughout the tested range, its coordination mechanisms may benefit from further optimization for very large deployments.
Importantly, DFUSE achieves this superior performance while maintaining strong consistency guarantees across all nodes and scales. The linear scaling pattern observed up to 16 nodes indicates that DFUSE’s architecture successfully addresses the scalability requirements of modern cloud environments without compromising on consistency semantics.
These results validate DFUSE’s design approach of offloading lease management to the kernel driver, demonstrating that it effectively balances performance, consistency, and scalability for distributed file system deployments of varying sizes.
@@ 7. Related Work
DFUSE builds upon and extends prior work in distributed file systems, FUSE optimizations, and consistency mechanisms in distributed file systems. This section contextualizes our contributions within the broader research landscape.
@@ 7.1. FUSE-based Distributed File Systems
Unlike these systems, DFUSE uniquely bridges the consistency-performance gap by enabling write-back caching with strong consistency through kernel-level lease management, avoiding the traditional tradeoff between performance and correctness.
@@ 7.2. FUSE Performance Optimizations
The Linux kernel has incorporated various FUSE optimizations, including the ability to splice read and write operations that move data between kernel buffers without copying to userspace(noa, 2010), significantly reducing memory copy overhead. Recent io_uring-based FUSE implementations(noa, 2024a) further improve performance by avoiding core switching and using shared memory for data transfers. These approaches collectively address the fundamental bottlenecks in FUSE: frequent context switches, memory copies, and lock contention.
These approaches primarily optimize local FUSE performance, whereas DFUSE addresses the fundamental distributed consistency challenge while leveraging these performance improvements. By offloading lease management to the kernel driver, DFUSE reduces the performance impact of maintaining strong consistency in distributed environments.
7.3. Consistency Mechanisms in Distributed File System
Distributed file systems employ various mechanisms to maintain consistency. Lease-based approaches, as pioneered by Gray and Cheriton(Gray and Cheriton, 1989), provide time-bound permissions that balance performance with consistency. GPFS (now Spectrum Scale)(Schmuck and Haskin, 2002) implements distributed locking for strong consistency but requires specialized hardware for optimal performance. Lustre(Schwan et al., 2003) employs a distributed lock manager to coordinate file access across nodes, though with different architectural tradeoffs than DFUSE.
DFUSE’s innovation lies in embedding lease management directly in the kernel’s FUSE driver while maintaining userspace flexibility. This approach differs fundamentally from previous systems that either implement consistency mechanisms entirely in userspace (sacrificing performance) or require extensive kernel modifications (sacrificing flexibility and isolation).
@@ 8. Conclusions
This paper presents DFUSE, a FUSE-based distributed file system that reconciles write-back kernel page caching with strong consistency for cloud environments. DFUSE eliminates the traditional trade-off between performance and correctness by offloading distributed lease management to the kernel driver, enabling coordinated access to cached data across nodes while retaining FUSE’s userspace flexibility. Our experiments demonstrate DFUSE’s efficacy: it achieves up to 68.0% higher throughput and 40.4% lower latency than write-through baselines for write-heavy workloads and near-linear scalability to 16 nodes while maintaining consistency under contention. Application benchmarks show 36.2% throughput improvements and 30.0% latency reduction in contended Fileserver workloads, validating its practicality for real-world deployments. By embedding lease coordination in the kernel, DFUSE demonstrates that distributed file systems can deliver both high performance and strong consistency, advancing the viability of FUSE-based solutions for cloud-scale workloads.
9. Acknowledgments
We thank the anonymous reviewers for their insightful feedback and helpful suggestions. This research was supported by NSF awards #2106530 and #2143868.